OpenStudy (anonymous):

Over two summers, Ray saved $800.00 and $600.00. The polynomial represents his savings at the beginning of the third year, where x is the growth factor. (The interest rate r is x – 1.) What is the interest rate he needs to save $1,650.00 after the third summer? 800x2 + 600x

OpenStudy (anonymous):

So, 800x^2 + 600x = 1650 From this you get 800x^2 + 600x - 1650 = 0 Now simplify this equation to get 16x^2 + 12x - 33 = 0 Plug this into the quadratic formula to get x.

OpenStudy (anonymous):

You are looking at about .11%

OpenStudy (anonymous):

or .1% if you have to round it.